OpenStudy (anonymous):

Please help! 1. Find k so that the line through (7, 2k) and (4, -3) is parallel to the line through (1,k+1) and (3,5).

OpenStudy (anonymous):

your question is incomplete.

OpenStudy (anonymous):

Thanks for catching that- I updated the correct question.

OpenStudy (anonymous):

slopes are same. \[slope=\frac{ y2-y1 }{ x2-x1 }\] equate the two slopes and find the value of k

OpenStudy (anonymous):

This is what I got: 1.) m= -3-2k/4-7=-3-2k/-3=-2k m=5-k+1/3-1=4-k/2=-k+4/2=-k What would I do to find the value of k? Would it be -2k=-1k?

myininaya (myininaya):

You simplify each slope incorrectly.

OpenStudy (anonymous):

Would the correct answer be k=6/7?

myininaya (myininaya):

\[m_1=\frac{-3-2k}{4-7}=\frac{-3-2k}{-3}\] \[m_2=\frac{5-(k+1)}{3-1}=\frac{5-k-1}{2}\]

OpenStudy (anonymous):

Thank you!!

myininaya (myininaya):

Well you still to find for what value of k m1=m2.

myininaya (myininaya):

And yes k=6/7. Good.

OpenStudy (anonymous):

So that is the final step?

myininaya (myininaya):

The last step is to state what k is.